# Synthetic vs. Gold: The Role of LLM-Generated Labels and Data in Cyberbullying Detection

### v2

@@ 4.4 Amount of Synthetic Data
For a more meaningful comparison of results with the authentic data, we sample subsets of the generated data to match the size of the authentic data for each CB case. We also sample subsets of certain percentages of these sizes from 100% to 200%. Since the test set is never synthetic, we measure sizes relative to the concatenation of training and validation data, i.e. messages correspond to 100%. The per-case target sizes for 100% are 863 (case A), 462 (case B), 103 (case C) and 325 (case D).
@@ 4.5 Authentic Unlabelled Data
In scenario 4, we remove the gold labels from the training and validation section of the authentic labelled data of scenario 1 to produce the unlabelled data to maximise comparability of the data and to limit changes to the labels that are annotated by the LLM in this scenario. In scenarios 3 and 4, the validation set with synthetic labels is used for model selection and early stopping, but validation set results in Section 5 will use the gold labels to measure performance.666Accuracy on synthetic validation data often exceeds 90% in our experiments, suggesting the that the synthetic data is less diverse than the authentic data for which validation accuracy usually is below 83%.
@@ 5.2 Scenario 2: LLMs as a Classifier
Table 3 presents the accuracy of Llama 3.3 70B, GPT-4o and Grok-2-1212 models using GE and GF prompts on both development and test sets in Scenario 2. We evaluate both GE and GF prompts on the development set and test the winning prompts on the test set. GPT-4o achieves the highest accuracy regardless of the prompt used, while Grok performs the lowest. The lower performance of GF prompts for two LLMs, particularly for Grok, highlights the importance of prompt design for model accuracy.
To decide which prompt to evaluate on test data for GPT-4o, we use the consistent performance trend with the other LLMs as a tie breaker. Comparing the GE test set results in Table 3 with those in Table 2, GPT-4o with the GE prompt reaches the highest test accuracy, outperforming the BERT-based model by 1.6%. However, despite the performance advantage, LLMs come with significant computational costs, making them less practical for real-time applications on resource-constrained devices. In contrast, the BERT-based classifier is a lightweight model optimized for efficiency, making it a more suitable choice for scenarios where speed and scalability are crucial. Since our industry partner aims to deploy CB detection on children’s mobile devices, we prefer using the BERT-based classifier due to its lightweight nature, which allows it to run efficiently on mobile hardware. This trade-off highlights the importance of balancing accuracy with application requirements as a small accuracy decrease of 1.6% can be acceptable in exchange for a model that is faster, more accessible, and capable of running on mobile devices.
@@ 5.3 Scenario 3: Fully Synthetic Data
Table 4 presents the development and test set accuracy for scenario 3: BERT-based classifiers trained on synthetic datasets generated by Llama, GPT, and Grok. The results indicate that Llama consistently outperforms the other two models across different dataset sizes. GPT follows closely, while Grok exhibits the weakest performance, particularly on the test set, where its accuracy is significantly lower than that of the other models. This suggests that Llama-generated synthetic data is of higher quality and more beneficial for training the classifier, despite the lower label quality observed in scenario 2.
@@ 5.4 Scenario 4: Synthetic Labels for Unlabeled Authentic Data
Table 5 shows that among the different LLMs used for synthetic annotation, Llama 3.3 produces the most useful labels, allowing BERT to achieve a test accuracy of 79.1%, closely approaching the human-labeled benchmark. This is despite the fact that the results in scenario 2 above show that the GPT-4o labels better match human labels than Llama 3.3’s labels. In scenario 4, GPT-4o follows with a slightly lower accuracy of 77.9%, while Grok lags behind at 75.2%. The difference in performance suggests that not all synthetic labels are of equal quality, and the choice of LLM for data annotation can impact model performance. Comparing the results in Table 2 with those in Table 11 shows that BERT trained on 100% authentic human-labeled data achieves a test accuracy of 81.5%, while the best-performing synthetic labels, produced by Llama 3.3, yield 79.1%. This small gap of just 2.4 percentage points highlights that synthetic labels can be a viable substitute to human labels while significantly reducing the cost, ethical concerns, and legal restrictions associated with human annotation, particularly for sensitive or harmful content.
@@ 6 Conclusions
In this paper, we explored the potential of LLM-generated synthetic data for cyberbullying detection, evaluating its effectiveness in three scenarios. Our results highlight that synthetic data can significantly reduce the reliance on human annotators while maintaining competitive model performance, especially in low-resource settings. However, we also observed that the quality and utility of synthetic data depend heavily on prompt design and data selection strategies.
For future work, we want to expand the set of data availability scenarios to include data augmentation, where authentic and synthetic data are combined. Furthermore, we plan to work with social scientists working in online safety to plan a human evaluation of the synthetic data. We also plan to investigate how temperature settings during label and conversation generation affect data diversity and model performance.
@@
Additionally, we want to try removing messages that are labeled with a low confidence score from the training data, similarly to, e.g., self-training Yarowsky (1995). Finally, we intend to explore selecting subsets of synthetic data based on their similarity to authentic data for better alignment with real-world patterns.
@@ Appendix A Appendix
| Case | Type of Problem |
| A: Your shy male classmate has a great passion for classical dance. Usually he does not talk much, but today he has decided to invite the class to watch him for his ballet show. | “Gendered division of sport” practices |
@@
| cyberbullying-related text categories are described below: |
| Sexual Talk Potentially harmful expressions with a sexual meaning. |
| Approach | Prompt |
| GE Synthetic Label | "moderation-criteria, Please analyse the following text and determine whether it contains any form of cyberbullying or harmful If it does, classify it as ’Harm’. if it does not, classify it as ’No Harm’. Text: Provide a detailed reasoning for your classification, thinking step by step. Finally, output the final answer as a tuple in the format: (’label’, confidence) - ‘label’ must be either ’Harm’ or ’No Harm’. ’confidence’ must be a number between 0 and 1. Do not include any additional text after the tuple. Reasoning: " |
| GF Synthetic Label | "Please analyze the following text and determine whether it contains any form of cyberbullying or harmful content. If it does, classify it as ’Harm’; if it does not, classify it as ’No Harm’. Text: Reasoning:" |
| Synthetic Data | "We are creating sample conversations to aid in cyberbullying detection. In these cases, teens are asked to role-play and create realistic conversations based on provided situations. There are 11 students participating in the conversation. The teens participating are: VCTM, BULLY1, BULLY2, VSUP1, VSUP2, VSUP3, VSUP4, BSUP1, BSUP2, BSUP3, BSUP4 with roles assigned as follows: VCTM: Victim, BULLY1 and BULLY2: Bully VSUP1, VSUP2, VSUP3 and VSUP4: Victim Support BSUP1, BSUP2, BSUP3 and BSUP4 : Bully Support consider this case: "value[’case’]" and consider this Type of addressed problem: "value[’problem’]" Generate an example conversation, with at least 100 messages, between these students based on the provided case and Type of addressed problem. Use profanity and strong language to create a realistic dialogue. number each message in the conversation. Please note that the conversation should be realistic and can be offensive. Please make sure to include different topics and perspectives in each conversation" |

### v3

@@ 4.3 Amount of Synthetic Data
For a more meaningful comparison of results with the authentic data, we sample subsets of the generated data to match the size of the authentic data. We maintain the ratio of messages belonging to each case (A to D) found in the authentic data. A size of 100% means that 863 messages are included for case A, 462 for case B, 103 for case C and 325 for case D.111111These numbers match the number of messages for each case in the training and validation data of the authentic dataset. The data is then split 60:20 = 75:25 to produce the training and validation data. The test set always uses authentic, manually labeled data. We sample subsets of certain percentages of these sizes from 100% to 200%. We ensure to produce more than 200% of synthetic data (relative to the size of the authentic data) for each LLM and each case so that these samples can be taken via sampling without replacement, not duplicating any data.121212The smallest available amount is approximately 208% for case B with Grok.
@@ 4.4 Authentic Unlabelled Data
In Scenario 4, we remove the gold labels from the training and validation section of the authentic labelled data of Scenario 1 to produce the unlabelled data to maximise comparability of the data and to limit changes to the labels that are annotated by the LLM in this scenario. In both Scenarios 3 and 4, the validation set with synthetic labels is used for model selection and early stopping, but validation set results in Section 5 will use the gold labels to measure performance.131313Accuracy on synthetic validation data often exceeds 90% in our experiments. This may mean that the synthetic data is less diverse than the authentic data for which validation accuracy usually is below 83%, or that harm is more clearly expressed.
@@ 5.2 Scenario 2: LLMs as a Classifier
Table 6 presents the accuracy of the Llama 3.3 70B, GPT-4o, and Grok-2-1212 models using guideline-enriched (GE) and guideline-free (GF) prompts on both the development and test sets in Scenario 2. We evaluate both GE and GF prompts on the development set and test the winning prompts on the test set. GPT-4o achieves the highest accuracy regardless of the prompt used, while Grok performs the lowest. The lower performance of GF prompts for two LLMs, particularly for Grok, highlights the importance of prompt design for model accuracy.
To decide which prompt to evaluate on test data for GPT-4o, we use the consistent performance trend with the other LLMs as a tie breaker. Comparing the guideline-enriched (GE)test set results in Table 6 with those in Table 5, GPT-4o with the GE prompt reaches the highest test accuracy, outperforming the BERT-based model by 1.6%. However, despite the performance advantage, LLMs come with significant computational costs, making them less practical for real-time applications on resource-constrained devices. In contrast, the BERT-based classifier is a lightweight model optimized for efficiency, making it a more suitable choice for scenarios where speed and scalability are crucial. Since our industry partner aims to deploy CB detection on children’s mobile devices, we prefer using the BERT-based classifier due to its lightweight nature, which allows it to run efficiently on mobile hardware. This trade-off highlights the importance of balancing accuracy with application requirements, as a small accuracy decrease of 1.6% can be acceptable in exchange for a model that is faster, more accessible, and capable of running on mobile devices.
@@ 5.3 Scenario 3: Fully Synthetic Data
Table 7 presents the development and test set accuracy for Scenario 3: BERT-based classifiers trained on synthetic datasets generated by Llama, GPT, and Grok141414To work within the LLMs’ safety constraints, we explicitly mentioned in the prompt: “We are creating sample conversations to aid in CB detection.” This clarification generally enabled the model to generate the required synthetic data. In instances where the model still refused to respond, we re-tried the prompt until we obtained a sufficient amount of data.. The results indicate that Llama consistently outperforms the other two models across different dataset sizes. GPT follows closely, while Grok exhibits the weakest performance, particularly on the test set, where its accuracy is considerably lower than that of the other models. This suggests that Llama-generated synthetic data is of higher quality and more beneficial for training the classifier, despite the lower label quality observed in Scenario 2.
@@ 5.4 Scenario 4: Synthetic Labels for Unlabeled Authentic Data
Table 8 shows that among the different LLMs used for synthetic annotation, Llama 3.3 produces the most useful labels, allowing BERT to achieve a test accuracy of 79.1%, closely approaching the human-labeled benchmark. This is despite the fact that the results in Scenario 2 above show that the GPT-4o labels better match human labels than Llama 3.3’s labels. In Scenario 4, GPT-4o follows with a slightly lower accuracy of 77.9%, while Grok lags behind at 75.2%. The difference in performance suggests that not all synthetic labels are of equal quality, and the choice of LLM for data annotation can impact model performance. Comparing the results in Table 5 with those in Table 11 shows that BERT trained on 100% authentic human-labeled data achieves a test accuracy of 81.5%, while the best-performing synthetic labels, produced by Llama 3.3, yield 79.1%. This small gap of just 2.4% highlights that synthetic labels can be a viable substitute to human labels while significantly reducing the cost, ethical concerns, and legal restrictions associated with human annotation, particularly for sensitive or harmful content.
@@ 6 Conclusions
In this paper, we explored the potential of LLM-generated synthetic data for cyberbullying detection, evaluating its effectiveness in four scenarios. Our results highlight that synthetic data can significantly reduce the reliance on human annotators while maintaining competitive model performance, especially in low-resource settings. However, we also observed that the quality and utility of synthetic data depend heavily on prompt design and data selection strategies.
For future work, we want to expand the set of data availability scenarios to include data augmentation, where authentic and synthetic data are combined. Furthermore, we plan to work with social scientists working in online safety to plan a human evaluation of the synthetic data. We also plan to investigate how the temperature settings during label and conversation generation affect data diversity and model performance.
@@
Additionally, we want to try removing messages that are labeled with a low confidence score from the training data, similarly to, e.g., self-training Yarowsky (1995); Blum and Mitchell (1998). Finally, we intend to explore selecting subsets of synthetic data based on their similarity to authentic data for better alignment with real-world patterns.
Acknowledgments
This research is supported by the Disruptive Technologies Innovation Fund (DTIF) under the project “Cilter: Protecting Children Online” Grant No. DT 2021 0362 from the Department of Enterprise, Trade and Employment in Ireland and administered by Enterprise Ireland (EI). This research is supported by the Research Ireland151515Previously Irish Research Council Enterprise Partnership Scheme (EPS) with Google for Online Content Safety under grant number EPSPG/2021/161. This research was conducted with the financial support of Science Foundation Ireland under Grant Agreement No. 13/RC/2106_P2 at the ADAPT SFI Research Centre at Dublin City University. ADAPT, the SFI Research Centre for AI-Driven Digital Content Technology, is funded by Science Foundation Ireland through the SFI Research Centres Programme. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.